# Flowception: Temporally Expansive Flow Matching for Video Generation

### v1

@@ Abstract
Figure 1: Examples of image-to-video (I2V) generation and video interpolation with Flowception . Input frames marked by dashed boundaries. Flowception enables variable-length non-autoregressive generation by learning to both denoise and insert frames in any order.
@@ 1 Introduction
Recent advances in diffusion and flow-matching have unlocked high-fidelity image generation Esser et al. (2024); Labs (2024) and are rapidly migrating to video. Current video generation models typically adopt one of two paradigms: full-sequence generation denoises all frames jointly with full attention Wan et al. (2025); HaCohen et al. (2024); Zheng et al. (2024), while temporal autoregressive (AR) generation produces frames (or blocks of frames) sequentially in a left-to-right order Bruce et al. (2024); Song et al. (2025). Full-sequence models benefit from bidirectional attention, enabling the model to correct errors during de-noising and achieving superior generation quality. However, parallel denoising the entire sequence prevents real-time streaming, as frames cannot be returned until fully denoised. Moreover, full-sequence models use a fixed generation length and incur a quadratic attention cost in the number of frames, limiting long-term generation.
@@
In contrast, AR methods enable streaming generation by making each frame immutable once generated, allowing immediate display to users while subsequent frames attend to prior outputs. However, these AR approaches suffer from critical exposure bias Zhang and Agrawala (2025); Bengio et al. (2015): training uses ground-truth frames as context, while inference conditions on the model’s own imperfect generations. This train-test mismatch prevents the model from learning to recover from its mistakes, causing minor artifacts to cascade as errors accumulate across frames, rapidly degrading video quality. Moreover, to enable KV caching, without which sampling from AR models becomes prohibitively expensive, AR methods typically use a causal attention mask which limits the expressiveness of the models. These trade-offs hinder long-term and efficient video synthesis.
@@
We conduct extensive experiments with three datasets, Tai-Chi-HD, RealEstate10K and Kinetics 600, and consider both image-to-video generation and video interpolation tasks. We find that, beyond being a more flexible framework, Flowception leads to comparable or better performance than full-sequence and autoregressive paradigms under the same compute budget, with consistent improvements in terms of FVD and VBench metrics.
@@
    We introduce Flowception , a theoretically grounded video generation framework, which couples frame insertions with continuous flow matching in a unified model.
@@
    We present an efficiency analysis showing an average of $3\times$ FLOPs reduction during training and ($1.5\times$ during sampling) w.r.t. full-sequence model.
@@ 2 Related work
Diffusion & Flows.  Diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Dhariwal and Nichol, 2021) and flow models (Lipman et al., 2023; Ma et al., 2024; Esser et al., 2024) are a class of generative models that operate by learning to model the infinitesimal transitions of the probability path going from source (noise distribution) to target (data manifold). They are considered state-of-the-art generative modeling tools for modalities ranging from images (Rombach et al., 2022; Podell et al., 2024; Chen et al., 2024; Esser et al., 2024), audio (Wang et al., 2023; Levy et al., 2023), and video (Jin et al., 2025; Polyak et al., 2025). These models generate data by starting with a simple prior such as a unit Gaussian, and iteratively refining samples to recover samples from the data distribution model. The simple training recipe underlying these models proved very effective through scalability and stable training Berrada et al. (2024).
@@
Video generation.  The success achieved by scaling text-to-image models Podell et al. (2024); Esser et al. (2024); Labs (2024) paved the way towards applying the same paradigms to video generation. Such works fall broadly into two categories. The first line of work Zheng et al. (2024); Peng et al. (2025); Wan et al. (2025) follows the full-sequence paradigm where the video is treated as one large tensor and all frames are denoised simultaneously. However, this results in prohibitively large attention computations, making the development of such methods prohibitive for long sequences.
@@
Another line of work Deng et al. (2024); Chen et al. (2025) takes an autoregressive approach to iteratively predict the next frame (or block of frames) conditioned on previous frames. This approach is known to have difficulty generating long videos, as sampling fixes early frames once they are generated and later frames are conditioned on (imperfectly) generated earlier frames (unlike in training, where conditioning is always based on ground-truth frames), leading to error accumulation that harms long-term generation. To resolve this, some methods rely on partial noising of the context Song et al. (2025), but this prevents KV caching, making sampling prohibitively expensive. Even without KV caching, our Flowception remains significantly faster than these autoregressive methods due to its interleaved insertion and denoising mechanism, whereas autoregressive generation is fundamentally constrained by sequential frame-by-frame generation.
@@
Another family of approaches factorizes long video generation into sparse keyframe synthesis followed by video interpolation. MovieDreamer Zhao et al. (2024) first autoregressively predicts a sequence of visual keyframes and then applies an image-to-video model to synthesize short clips around each keyframe, using CLIP features of the anchor frame rather than the last generated frame to mitigate drift across clips. ART-V Weng et al. (2024) instead performs frame-wise autoregressive generation in a text-image-to-video (TI2V) setting: each new frame is produced by a diffusion model conditioned on both the input keyframe and the previously generated frames, with a masked diffusion mechanism that selectively reuses information from the given frame to reduce appearance drift over long horizons.
@@
Several other directions have been explored to reduce complexity. These include the use of highly compressive tokenizers HaCohen et al. (2024); NVIDIA et al. (2025), hierarchical or packed representations that compress remote frames before attending to them Jin et al. (2025); Zhang and Agrawala (2025), and architectural changes that distribute or route attention over multiple context sources Fan et al. (2025); Cai et al. (2025); Song et al. (2025); Zhao et al. (2024). However, these approaches still depend on design choices about which context sources or time ranges to emphasize, rather than fully learning relevance end-to-end.
@@
To enable multiple use cases, such as image-to-video and video-to-video, some works Wan et al. (2025); HaCohen et al. (2024) concatenate a frame-wise mask to the input of the model, which allows to condition on clean frames at various positions in the video. Similarly, Flowception is capable of performing various tasks by conditioning on multiple input frames, and optionally constraining insertions to be to the right of some of those.
@@ 3 Flowception
We propose a natural combination of the continuous Flow Matching (Lipman et al., 2023) and the discrete Edit Flow (Havasi et al., 2025) frameworks, with the goal of flexible video generation where existing frames are simultaneously denoised and new frames are inserted during the generation process. This is our Flowception model. We first describe the parameterization and the generation procedure for this model, then derive a training scheme that adheres to the generation procedure. We will present notation loosely with high-level intuitive explanations; full derivations are in the supplementary material.
@@ 3.2 Generation procedure
Global time value.  The stopping criterion for the flow is when $t^{i}=1$ for all frames, i.e. we no longer modify any frame. To impose a stopping criterion for the insertions, we introduce a global time value $t_{g}$ which starts at $t_{g}=0$ and frame insertions are only allowed while $t_{g}<1$. Implementation-wise, we don’t feed $t_{g}$ explicitly into the model, and is only tracked for sampling. This closely follows the design choice of OneFlow Nguyen et al. (2025) where insertion rates are not conditioned on time. Accordingly, with a constant step size during inference, the maximum number of iterations is twice the amount needed for $t_{g}$ to arrive to $1$.
@@
Transport step.  Given a sequence $X$ with per-frame times $t$, each transport step performs the following two operation simulatenuosly at all positions $i\in\{1,\dots,\ell(X)\}$:
@@
The ratio $\tfrac{\dot{\kappa}(t_{g})}{1-\kappa(t_{g})}$ ensures insertions occur in aligment with the distribution of visible frames imposed by the scheduler.
@@ 3.3 Training procedure
Secondly, while $\tau_{g}\leq 1$, insertions occur according to the scheduler $\kappa(\cdot)$. For the linear scheduler—for the general case, please see the detailed derivations in the supplementary material—new frames are inserted uniformly across the time interval $t_{g}\in[0,1]$. Therefore, the time delay between the global extended time and per-frame extended time values follows:
@@
where $t_{i}=\text{clip}(\tau_{i})$ computes the real per-frame times. Note that $\tau_{i}\in[-1,2]$ according to (7), When $\tau_{i}<0$, the frame is yet to be inserted and thus in a “deleted” state.
@@
Insertion loss.  At each position $i$ of the noisy sequence $X$, we denote $k^{i}$ as the number of missing frames to the right, between positions $i$ and $i+1$. Following OneFlow Nguyen et al. (2025), the loss for training the insertion output $\lambda_{i}$, which predicts the number of missing frames, is given by the negative log-likelihood of the Poisson distribution
@@ 3.4 Practical considerations & implications
Architecture with per-frame time conditioning.  Differently from other frameworks, our model has two prediction heads: the dense velocity prediction head and the insertion rate prediction head. In order to enable different timesteps per frame, we change the AdaLN Peebles and Xie (2023) modulation to operate on a per-frame basis. To distinguish between noisy and input/conditioning frames, we inflate the channel dimension of the inputs so that the first $C$ channels contain the noisy frames while the second group contains the input/conditioning frames (or zero padding). Finally, for the rate prediction head, a learnable token is concatenated to each frame’s tokens before being projected with a simple MLP and an exponential activation.
@@ 4 Experiments
Figure 5: Image-to-video generation using model trained on Tai-Chi-HD. We show two examples (one per row) of generated video frames along the insertion time of each frame. In general, frames inserted early, with $t\approx 0$, define the movement dynamics (large changes w.r.t. the context frame), while later frames smoothly interpolate motion, resulting in smaller changes w.r.t. to neighboring frames.
@@ 4.1 Experimental setup
Model architecture.  Our architecture builds on the popular DiT diffusion transformer architecture Peebles and Xie (2023), with minimal changes to enable insertion rate predictions. We equip each attention block with an additional learnable token which is replicated for each frame and takes part in the attention computation, ultimately each of these tokens maps to the non-negative rate prediction using an exponential activation. Our baseline model is made of 38 attention blocks with a hidden dimension of 1536, each attention head has a dimension of 64, resulting in 24 total heads, for a model with approximately 2.1B learnable parameters. Positional embeddings follow the VideoROPE method Wei et al. (2025). We enable full bidirectional attention between frames and optionally concatenated text tokens in an MMDiT Esser et al. (2024) fashion.
Datasets.  We evaluate our model on three different datasets. We used two narrow-domain datasets: Tai-Chi-HD Siarohin et al. (2019) and RealEstate10K Zhou et al. (2018), as well as the class-structured Kinetics 600 dataset Carreira et al. (2018). Unless specified otherwise, we train our models at $256$ resolution and generate up to 145 frames at 16 FPS. Furthermore, we trained a general domain model using a proprietary dataset of around 20M videos ranging approximately from 20 to 240 seconds. We show results for this model in the supplementary material. We use the LTX autoencoder HaCohen et al. (2024), which has a spatial downsampling factor of 32 and temporal downsampling factor of 8. To correctly support variable length video decoding, we modify the decoder architecture to propagate a frame validity mask which is used to mask out invalid frames in the hidden layers, preventing border effects where padding frames leak into the video. When prompting models with one or more video frames to generate a video, we choose these reference frames from the validation set.
Metrics.  To evaluate the performance of our models, we rely on the standard FVD metric Unterthiner et al. (2019), computed with respect to a subset of 5k videos from the training set, as well as selected metrics from VBench Huang et al. (2024). Specifically, we report imaging and aesthetic quality, background and subject consistency, motion smoothness and dynamic degree.
@@ 4.2 Main results
Figure 6: Video interpolation results obtained with with Flowception trained on the Kinetics-600 dataset. Context frames are highlighted with dashed lines, insertion times of other frames are marked using a color map and printed on each frame.
Video interpolation.  Another interesting task directly supported by Flowception is video interpolation, where rather than a single starting frame, we provide an ordered set of context frames as conditioning, and let the model interpolate the video freely between these. Differently from other frameworks Wan et al. (2025); HaCohen et al. (2024), Flowception does not need to be given the number of frames between successive conditioning frames, but it rather inserts frames as it sees fit, resulting in a more flexible interpolation capability. We illustrate this use-case in Figure˜6. We observe that the model is able to adapt the number of insertions to generate coherent scenes with smooth transitions. For instance, in the second row, no frame is inserted after the penultimate context frame in order to preserve motion continuity.

### v2

@@ Abstract
Figure 1: Example generations for Flowception under different modalities. Context frames marked by dashed boundaries. Flowception enables variable-length non-autoregressive generation by learning to both denoise and insert frames in any order as sampling progresses (captions in the supplementary).
@@ 1 Introduction
Recent advances in diffusion and flow-matching have unlocked high-fidelity image generation (Esser et al., 2024; Labs, 2024) and are rapidly migrating to video. Current video generation models typically adopt one of two paradigms: full-sequence generation denoises all frames jointly with full attention (Wan et al., 2025; HaCohen et al., 2024; Zheng et al., 2024), while temporal autoregressive (AR) generation produces frames (or blocks of frames) sequentially in a left-to-right order (Bruce et al., 2024; Song et al., 2025). Full-sequence models benefit from bidirectional attention, enabling the model to correct errors during denoising and achieving superior generation quality. However, parallel denoising of the entire sequence prevents real-time streaming, as frames cannot be returned until fully denoised. Moreover, full-sequence models use a fixed generation length and incur a quadratic attention cost in the number of frames, limiting long-term generation.
@@
In contrast, AR methods enable streaming generation by making each frame immutable once generated, allowing immediate display to users while subsequent frames attend to prior outputs. However, these AR approaches suffer from critical exposure bias (Zhang and Agrawala, 2025; Bengio et al., 2015): training uses ground-truth frames as context, while inference conditions on the model’s own imperfect generations. This train-test mismatch prevents the model from learning to recover from its mistakes, causing minor artifacts to cascade as errors accumulate across frames, rapidly degrading video quality. Moreover, to enable KV caching, without which sampling from AR models becomes prohibitively expensive, AR methods typically use a causal attention mask which limits the expressiveness of the models. These trade-offs hinder long-term and efficient video synthesis.
@@
We conduct extensive experiments with three datasets, Tai-Chi-HD, RealEstate10K and Kinetics 600, and consider (text/image)-to-video generation and video interpolation tasks. We find that, beyond being a more flexible framework, Flowception leads to comparable or better performance than full-sequence and autoregressive paradigms under the same compute budget, with consistent improvements in terms of FVD and VBench metrics.
@@
    We introduce Flowception, a theoretically grounded video generation framework, which couples learned frame insertions with continuous flow matching in a unified model.
@@
    We present an efficiency analysis showing an average of $3\times$ FLOPs reduction during training, and $1.5\times$ during sampling, w.r.t. full-sequence model.
@@ 2 Related work
Diffusion & Flows.  Diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Dhariwal and Nichol, 2021) and flow models (Lipman et al., 2023; Ma et al., 2024; Esser et al., 2024) are a class of generative models that operate by learning to model the infinitesimal transitions of the probability path going from source (noise distribution) to target (data manifold). They are considered state-of-the-art generative modeling tools for modalities ranging from images (Rombach et al., 2022; Podell et al., 2024; Chen et al., 2024; Esser et al., 2024), audio (Wang et al., 2023; Levy et al., 2023), and video (Jin et al., 2025; Polyak et al., 2025). These models generate data by starting with a simple prior such as a unit Gaussian, and iteratively refining samples to recover samples from the data distribution model. The simple training recipe underlying these models proved very effective through scalability and stable training (Berrada et al., 2024).
@@
Video generation.  The success achieved by scaling text-to-image models (Podell et al., 2024; Esser et al., 2024; Labs, 2024) paved the way towards applying the same paradigms to video generation. Such works fall broadly into two categories. The first line of work (Zheng et al., 2024; Peng et al., 2025; Wan et al., 2025) follows the full-sequence paradigm where the video is treated as one large tensor and all frames are denoised simultaneously. However, this results in prohibitively large attention computations, hindering the development of such methods for long sequences.
@@
Another line of work (Deng et al., 2025; Chen et al., 2025) takes an autoregressive approach to iteratively predict the next frame (or block of frames) conditioned on previous frames. This approach is known to have difficulty generating long videos, as sampling fixes early frames once they are generated and later frames are conditioned on (imperfectly) generated earlier frames (unlike in training, where conditioning is always based on ground-truth frames), leading to error accumulation that harms long-term generation. To resolve this, some methods rely on partial noising of the context (Song et al., 2025), but this prevents KV caching, making sampling prohibitively expensive. Even without KV caching, our Flowception remains significantly faster than these autoregressive methods due to its interleaved insertion and denoising mechanism, whereas autoregressive generation is fundamentally constrained by sequential frame-by-frame generation.
@@
Another family of approaches factorizes long video generation into sparse keyframe synthesis followed by video interpolation. MovieDreamer (Zhao et al., 2025) first autoregressively predicts a sequence of visual keyframes and then applies an image-to-video model to synthesize short clips around each keyframe, using CLIP features of the anchor frame rather than the last generated frame to mitigate drift across clips. ART-V (Weng et al., 2024) instead performs frame-wise autoregressive generation in a text-image-to-video (TI2V) setting: each new frame is produced by a diffusion model conditioned on both the input keyframe and the previously generated frames, with a masked diffusion mechanism that selectively reuses information from the given frame to reduce appearance drift over long horizons.
@@
Several other directions have been explored to reduce complexity. These include the use of highly compressive tokenizers (HaCohen et al., 2024; NVIDIA et al., 2025), hierarchical or packed representations that compress remote frames before attending to them (Jin et al., 2025; Zhang and Agrawala, 2025), and architectural changes that distribute or route attention over multiple context sources (Fan et al., 2025; Cai et al., 2026; Song et al., 2025; Zhao et al., 2025). However, these approaches still depend on design choices about which context sources or time ranges to emphasize, rather than fully learning relevance end-to-end.
More related to our work, several works explored an autoregressive approach, but denoise multiple frames simultaneously, with later frames being noisier than earlier frames (Teng et al., 2025; Xie et al., 2025). Differently from Flowception, however, these approaches maintain a strict left-to-right frame insertion pattern, preventing coordination between remote frames early in the denoising process as Flowception allows it.
@@
To enable multiple use cases, such as image-to-video and video-to-video, some works (Wan et al., 2025; HaCohen et al., 2024) concatenate a frame-wise mask to the input of the model, which allows to condition on clean frames at various positions in the video. Similarly, Flowception is capable of performing various tasks by conditioning on multiple input frames, and optionally constraining insertions to be to the right of some of those.
@@ 3 Flowception
We propose Flowception, a natural combination of the continuous Flow Matching (Lipman et al., 2023) and the discrete Edit Flow (Havasi et al., 2025) frameworks, with the goal of flexible video generation where existing frames are simultaneously denoised and new frames are inserted during the generation process. We first describe the parameterization and the generation procedure for this model, then derive a training scheme that adheres to the generation procedure. We will present notation loosely with high-level intuitive explanations; full derivations are in the supplementary material.
@@ 3.2 Generation procedure
Global time value.  The stopping criterion for the flow is when $t^{i}=1$ for all frames, i.e. we no longer modify any frame. To impose a stopping criterion for the insertions, we introduce a global time value $t_{g}$ which starts at $t_{g}=0$ and frame insertions are only allowed while $t_{g}<1$. Implementation-wise, we don’t feed $t_{g}$ explicitly into the model, and is only tracked for sampling. This closely follows the design choice of OneFlow (Nguyen et al., 2025) where insertion rates are not conditioned on time. Accordingly, with a constant step size during inference, the maximum number of iterations is twice the amount needed for $t_{g}$ to arrive to $1$.
@@
Transport step.  Given a sequence $X$ with per-frame times $t$, each transport step performs the following two operation simultaneously at all positions $i\in\{1,\dots,\ell(X)\}$:
@@
The ratio $\tfrac{\dot{\kappa}(t_{g})}{1-\kappa(t_{g})}$ ensures insertions occur in alignment with the distribution of visible frames imposed by the scheduler.
@@ 3.3 Training procedure
Secondly, while $\tau_{g}\leq 1$, insertions occur according to the scheduler $\kappa(\cdot)$. For the linear scheduler new frames are inserted uniformly across the time interval $t_{g}\in[0,1]$. For the general case, please see the detailed derivations in the supplementary material. Therefore, the time delay between the global extended time and per-frame extended time values follows:
@@
where $t_{i}=\text{clip}(\tau_{i})$ computes the real per-frame times. Note that $\tau_{i}\in[-1,2]$ according to (7). When $\tau_{i}<0$, the frame is yet to be inserted and thus in a “deleted” state.
@@
Insertion loss.  At each position $i$ of the noisy sequence $X$, we denote $k^{i}$ as the number of missing frames to the right, between positions $i$ and $i+1$. Following OneFlow (Nguyen et al., 2025), the loss for training the insertion output $\lambda_{i}$, which predicts the number of missing frames, is given by the negative log-likelihood of the Poisson distribution
@@ 3.4 Practical considerations & implications
Architecture with per-frame time conditioning.  Differently from other frameworks, our model has two prediction heads: the dense velocity prediction head and the insertion rate prediction head. In order to enable different timesteps per frame, we change the AdaLN (Peebles and Xie, 2023) modulation to operate on a per-frame basis. To distinguish between noisy and input/conditioning frames, we inflate the channel dimension of the inputs so that the first $C$ channels contain the noisy frames while the second group contains the input/conditioning frames (or zero padding). Finally, for the rate prediction head, a learnable token is concatenated to each frame’s tokens before being projected with a simple MLP and an exponential activation.
@@ 4 Experiments
Figure 5: Image-to-video generation using model trained on Tai-Chi-HD. We show two examples (one per row) of generated video frames along the insertion time of each frame. In general, frames inserted early, with $t\approx 0$, define the movement dynamics (large changes w.r.t. the context frame), while later frames smoothly interpolate motion, resulting in smaller changes w.r.t. neighboring frames.
@@ 4.1 Experimental setup
Model architecture.  Our architecture builds on the popular DiT diffusion transformer architecture (Peebles and Xie, 2023), with minimal changes to enable insertion rate predictions. We equip each attention block with an additional learnable token which is replicated for each frame and takes part in the attention computation, ultimately each of these tokens maps to the non-negative rate prediction using an exponential activation. Our baseline model is made of 38 attention blocks with a hidden dimension of 1536, each attention head has a dimension of 64, resulting in 24 total heads, for a model with approximately 2.1B learnable parameters. Positional embeddings follow the VideoROPE method (Wei et al., 2025). We enable full bidirectional attention between frames and optionally concatenated text tokens in an MMDiT (Esser et al., 2024) fashion. For T2V results, we finetune an open-source LTX-2b model (0.9.5) (HaCohen et al., 2024) with our framework after conducting the architecture changes necessary for Flowception.
Datasets.  We evaluate our model on three different datasets. We used two narrow-domain datasets: Tai-Chi-HD (Siarohin et al., 2019) and RealEstate10K (Zhou et al., 2018), as well as the class-structured Kinetics 600 dataset (Carreira et al., 2018). Unless specified otherwise, we train our models at $256$ resolution and generate up to 145 frames at 16 FPS. Furthermore, we trained a general domain model using a proprietary dataset of around 2M videos ranging approximately from 20 to 240 seconds. We show results for this model in the supplementary material. We use the LTX autoencoder (HaCohen et al., 2024), which has a spatial downsampling factor of 32 and temporal downsampling factor of 8. To correctly support variable length video decoding, we modify the decoder architecture to propagate a frame validity mask which is used to mask out invalid frames in the hidden layers, preventing border effects where padding frames leak into the video. When prompting models with one or more video frames to generate a video, we choose these reference frames from the validation set.
Metrics.  To evaluate the performance of our models, we rely on the standard FVD metric (Unterthiner et al., 2019), computed with respect to a subset of 5k videos from the training set, as well as selected metrics from VBench (Huang et al., 2024). Specifically, we report imaging and aesthetic quality, background and subject consistency, motion smoothness and dynamic degree.
@@ 4.2 Main results
Text-to-video (T2V) generation.  We train Flowception on the proprietary dataset of around 2M text-video pairs described in Section˜4.1. We use a pretrained LTX model for this experiment as stated earlier. Our results are reported in Table˜2, Flowception consistently improves perceptual quality and temporal coherence over the full-sequence paradigm, achieving higher imaging quality, background consistency, aesthetics, motion smoothness and subject consistency while dynamic degree is only slightly deteriorated. Qualitatively, Flowception produces sharper motion and fewer identity drifts over long time horizons (Fig. 1). We empirically observe similar trends to I2V where early frames define coarse motion while later inserted frames ensure smooth transitions between key frames.
[…]
LTX (original) 49.22 95.13 48.25 98.49 92.30 53.90
LTX (full-sequence) 47.96 93.82 47.28 98.07 89.47 53.71
LTX (Flowception) 51.37 94.95 49.56 98.14 91.40 52.43
Table 2: Text-to-video generation results. We adapt LTX-2b model and compare Flowception with full-sequence training.
Figure 6: Frame interpolation results obtained with Flowception trained on the Kinetics-600 dataset. Context frames are highlighted with dashed lines, insertion times of other frames are marked using a color map and printed on each frame.
Frame interpolation.  Another interesting task directly supported by Flowception is frame interpolation, where rather than a single starting frame, we provide an ordered set of context frames as conditioning, and let the model interpolate the video freely between these. Differently from other frameworks (Wan et al., 2025; HaCohen et al., 2024), Flowception does not need to be given the number of frames between successive conditioning frames, but it rather inserts frames as it sees fit, resulting in a more flexible interpolation capability. We illustrate this use-case in Figure˜6. We observe that the model is able to adapt the number of insertions to generate coherent scenes with smooth transitions. For instance, in the second row, no frame is inserted after the penultimate context frame in order to preserve motion continuity.
@@
Sampling Efficiency.  We compare wall-clock time for sampling the same LTX-2b model used for our T2V experiment. We track sampling time on a single H200 GPU with different numbers of sampling steps (NFEs). Flowception is consistently about 30% faster than the Full-Sequence baseline.
Figure 7: Sampling speed. We compare wall-clock time when sampling videos using either full-sequence or Flowception paradigms.